# zkLoRA: Fine-Tuning Large Language Models with Verifiable Security via Zero-Knowledge Proofs

### v1

@@ V-B Completeness
Statement: zkLoRA ensures completeness, meaning that a verifier will always accept proofs generated by an honest PPT prover who accurately performs all computations as specified in the zkLoRA protocol described in Protocol 2.
Analysis: We analyze the completeness of zkLoRA in its arithmetic computations and non-arithmetic computations, respectively.
  • Arithmetic Operations: zkLoRA proves the arithmetic computations in LoRA using the sumcheck protocol, which is known to have perfect completeness. This implies that when the prover is honest and performs all computations correctly, the verifier will accept the proof with probability exactly one.
  • Non-arithmetic Operations: zkLoRA proves the non-arithmetic computations in LoRA by employing lookup-based arguments. Completeness may fail if the random challenge from the verifier (e.g., $\beta$) coincides with a pole of the rational function used in the identity check of lookup-based arguments  [7]. However, the probability of such an event is bounded by $\frac{C}{|\mathbb{F}|}$, which is negligible for a sufficiently large field size.
In summary, zkLoRA ensures that the completeness error is negligible, guaranteeing that an honest prover who performs all computations correctly will have an overwhelming probability of convincing the verifier to accept the proof.
@@ V-C Zero Knowledge
Statement: zkLoRA ensures zero knowledge, meaning that the verifier learns nothing beyond the validity of the statement being proven. Formally, there exists a polynomial-time simulator $\mathcal{S}$ capable of generating transcripts indistinguishable from real interactions, even without access to the prover’s secret witness. The zero-knowledge error is negligible, bounded by $\mathsf{negl}(\lambda)$ under standard cryptographic assumptions.
Analysis: The zero-knowledge property of zkLoRA is achieved by incorporating the Hyrax-style polynomial commitments [30] into the sumcheck protocol [7], which provides computational hiding under the discrete logarithm assumption in the random oracle model. Specifically, zkLoRA ensures that the verifier learns nothing beyond the validity of the computation while keeping private: 1) the pre-trained model parameters, 2) the low-rank matrices and their updates, 3) the internal computation values in SwiGLU, Softmax, transpose, and element-wise operations during the forward and backward propagations, 4) the training data in the mini-batch.
To rigorously establish the zero-knowledge property of zkLoRA, we define two executions from the perspective of any PPT adversary $\mathcal{A}$: a real execution with an honest prover and an ideal execution with a simulator. The security objective is to guarantee that the adversary cannot distinguish these two executions.
Formally, zkLoRA is zero-knowledge if there exists a polynomial-time simulator $\mathcal{S}$ such that, for all PPT adversaries $\mathcal{A}$ and for all witness $w$ (containing pre-trained parameters, LoRA matrices, intermediate computation values, and training data), the following distributions are computationally indistinguishable: $\left|\Pr\left[\mathcal{A}(\mathrm{Real}(w,\mathrm{pp}))=1\right]-\Pr\left[\mathcal{A}(\mathrm{Ideal}(\mathcal{S},\mathrm{pp}))=1\right]\right|\leq\mathsf{negl}(\lambda)$, where $\mathrm{Real}(w,\mathrm{pp})$ and $\mathrm{Ideal}(\mathcal{S},\mathrm{pp})$ denote the real and ideal executions, respectively, and $\mathrm{pp}$ denotes public parameters. The definitions of $\mathrm{Real}$ and $\mathrm{Ideal}$ are given below.
$\mathrm{Real~Execution}~\mathrm{Real}(w,\mathrm{pp})$: 1. $com\leftarrow zkLoRA\text{-}Commit(w;pp)$, where $w$ includes the pre-trained parameters, the low rank matrices and their updates, the intermediate computation values in all operations, and the training batch. 2. $\pi\leftarrow zkLoRA\text{-}Prove(com;pp)$, using the zkMat, zkElementProd, zkTranspose, zkSwiGLU, and zkSoftmax protocols. 3. return $com,\pi$
@@
In summary, zkLoRA ensures that the zero-knowledge error is negligible, guaranteeing that the verifier learns nothing beyond the validity of the statement, while preserving the privacy of the prover’s secret witness.
Figure 2: The proving time (in seconds) for different models: (a) the forward propagation phase; (b) the backward propagation phase; (c) the parameter update phase, (d) the total.
@@ VI Experimental Evaluations
To validate the practical feasibility and performance of zkLoRA, we conducted comprehensive experiments on real-world large language models and datasets. This section outlines the experimental setup, implementation details, and evaluation results.
@@ VI-A Implementation Details
The implementation of zkLoRA adopts a hybrid architecture where CUDA components handle core cryptographic operations and zero-knowledge proof generation, while Python scripts provide high-level orchestration and model-specific interfaces. The CUDA modules implement elliptic curve arithmetic and pairing operations on the BLS12-381 curve [31], cryptographic commitments for model parameters, and GPU-accelerated zero-knowledge protocols for LoRA Fine-tuning on LLMs. The Python layer manages model downloading, file I/O utilities, and LLaMA-specific processing workflows. We evaluated zkLoRA for LoRA fine-tuning across six Transformer-based LLMs of varying scales, including LLaMA-3.2 [2] with 3B and 11B parameters, LLaMA-2 [2] with 7B and 13B parameters, and OPT [32] with 6.7B and 13B parameters, covering models up to 13 billion parameters. We trained the LLMs with the public C4 dataset of English-language text [30]. Our code is publicly available at https://github.com/liaoguofu/zkLoRA.
@@ VI-D1 Proving Time
We evaluated the prover runtime of zkLoRA across the six Transformer based LLMs. All measurements correspond to the the training process of a mini-batch consists of a single data sample. For each model, we measured the prover runtime of three distinct computational phases of LoRA: forward propagation, backward propagation, and parameter updates. As shown in Fig. 2, the total proving time ranges from 121.93 seconds (LLaMA-3.2-3B) to 249.38 seconds (OPT-13B), reflecting the increasing computational overhead as model size scales.
@@ VI-D2 Verification Time
We measured verification time for the six tested LLMs and present the results in Table I. Across all tested LLMs, verification remains highly efficient. As shown in Table I, the end-to-end verification time ranges from 1.87 seconds (LLaMA-3.2-3B) to 3.73 seconds (OPT-13B), consistently staying below 4 seconds even for models with over 13 billion parameters. These results demonstrate that the verification workload in zkLoRA is much more efficient than the proof generation and represents only a minor fraction of the total computational cost.
@@ VI-D3 Polynomial Commitment Cost
We measured the time costs and sizes of the Hyrax polynomial commitments used in zkLoRA. From the results provided in Table I, the generation of the commitments introduces a substantial time cost, particularly for large-scale models. The commitment generation time ranges from 156 seconds for LLaMA-3.2-3B to 554 seconds for OPT-13B. This step involves producing polynomial commitments for all relevant intermediate computations and dominates the wall-clock latency in our evaluation. The commitment sizes follow a similar trend, increasing from 135.59 MB to 232.67 MB, consistent with the growth in matrix and lookup table dimensionality.
@@ VI-D4 GPU Memory Usage
The total memory usage reported in Table I represents the cumulative GPU memory allocated across all Transformer layers during the forward propagation, backward propagation, and parameter update phases while executing zkLoRA. These memory usages range from approximately 2573 GB for LLaMA-3.2-3B to 3920 GB for LLaMA-2-11B. The computations for the forward propagation, backward propagation, and parameter update phases within the sub-layers of the Transformer layers are executed sequentially on the GPU. Consequently, the actual peak memory observed during zkLoRA execution remains below 80 GB. This allows zkLoRA to be efficiently processed on a single NVIDIA A100 GPU with 80 GB of device memory, ensuring practicality even for the largest model scale evaluated.
@@
Model Verification Time (s) Commitment Time (s) Commitment Size (MB) Total Memory Usage (GB) LLaMA-3.2 3B 1.87 156 135.59 2573 LLaMA-3.2 11B 2.99 299 228.31 3920 LLaMA-2 7B 2.23 232 182.84 2933 LLaMA-2 13B 3.49 304 224.56 3902 OPT 6.7B 1.95 285 165.48 3135 OPT 13B 3.73 554 232.67 3912
VII Related work
In this section, we survey verifiable machine learning from three perspectives: i) verifiable inference, which focuses on proving the correctness of a model’s forward propagation under zero knowledge; ii) verifiable training, which aims to authenticate gradient computation and parameter updates; and iii) verifiable testing, which certifies global performance metrics (e.g. accuracy or fairness) without revealing model or data.
VII-A Verifiable Inference
Early works like SafetyNets [4] used interactive proofs to verify DNN predictions without privacy protection. Subsequent zero-knowledge approaches like vCNN [33] introduced efficient convolution encodings for zk-SNARKs.
For Transformers, ZKML [6] provided the first end-to-end zero-knowledge proof for GPT-2 inference using Halo2 lookup arguments, though with substantial proving overhead. zkLLM [7] introduced GPU acceleration, reducing latency and scaling to 13B parameters.
Most recently, zkGPT [34] achieved sub-25 second proofs for GPT-2 (117M) through constraint fusion and circuit optimization, outperforming prior work by 185-279×.
Collectively, these efforts demonstrate that ZKPs can authenticate machine learning inference steps—the model owner (prover) can convince a client (verifier) that a given prediction was computed correctly by a neural network, without revealing inputs or weights. However, all of the above works target the inference procedure of fixed, pre-trained models. In contrast, our work focuses on proving the training (fine-tuning) process itself, which involves a series of parameter updates rather than a single forward-pass computation.
VII-B Verifiable Training
Verifying training is more demanding than inference due to gradient computation. Early approaches side-stepped this complexity by verifying only snapshots. VeriML [5] asks the service provider to commit to intermediate states and later supply ZK proofs for randomly sampled iterations, covering linear and shallow-net models but not deep networks. Distributed settings such as Drynx [35] and zkMLaaS [36] combine MPC and interactive proofs to check aggregated updates, again limited to low-complexity learners.
To support full deep neural network (DNN) training, Sun and Zhang introduced zkDL [29], which developed a custom zkReLU protocol to prove forward and backward propagation through ReLU activations. zkDL flattens the training process into a parallel circuit architecture (FAC4DNN), enabling the generation of a single-epoch proof for CNNs with tens of millions of parameters in under a minute.
Building on the zkPoT framework, Abbaszadeh [37] further optimized per-step GKR-based sumchecks and recursively aggregated the results across gradient steps, yielding concise proofs across multiple training iterations for moderate-sized models.
Beyond arithmetic correctness, Shamsabadi et al. [38] introduced Confidential-PROFITT, which provides zero-knowledge proofs of fair training for decision trees. Their protocol certifies that the trained model satisfies demographic parity constraints, without revealing model parameters or sensitive attributes, and completes in minutes.
Despite these advances, no prior work supports verifiable fine-tuning of large-scale Transformer models. In particular, LoRA [3], a widely adopted technique that adapts only low-rank matrices while freezing the base model, has not been addressed in prior ZKP systems. Our work is the first to construct a zero-knowledge proof system for LoRA fine-tuning, verifying that parameter updates are computed correctly on the specified training data while preserving confidentiality.
VII-C Verifiable Testing
Verifiable testing seeks to certify global properties—such as accuracy or fairness—of a trained model on an evaluation set, while hiding the model and inputs. The first dedicated ZKP for this task is zkDT [39], which proves both per-sample inference and dataset-wide accuracy for decision-tree models by embedding all test paths in a single Aurora proof; a 23-level tree evaluated on 5000 CIFAR samples yields a 250s proof of only 287 KB. Campanelli et al. extend this line with cq+/zkcq+ matrix-lookup arguments [40], letting the prover commit to the entire decision tree as a matrix and isolate only the rows reached by all test inputs. Their zero-knowledge matrix lookup reduces prover time to be independent of the tree size, substantially cutting the overhead of zkDT.
For convolutional networks, pvCNN [41] combines homomorphic encryption, Quadratic-Matrix-Program (QMP) circuits, and proof aggregation to batch-verify accuracy across many testers; its QMP encoding yields a 13.9× speed-up over traditional QAP-based zk-SNARKs for high-dimensional convolutions. ZEN [42] introduces a compiler that quantizes floating-point CNNs into R1CS with sign-bit grouping and SIM(D) stranded encoding, reducing constraints by up to 22× with no accuracy loss. zkCNN [43] further optimises two-dimensional convolution and FFT inside Groth16, achieving end-to-end proofs for VGG-16 in 88 s with 341 KB proofs and 59 ms verification.
Collectively, these systems demonstrate that ZKPs can certify model-wide statistics or properties without exposing model weights or evaluation data. However, they all presume access to a fixed, pre-trained model and do not address the fundamental challenge of verifying how the model was trained or fine-tuned. This limitation is critical in scenarios where the training process itself must be auditable and trustworthy. Our work addresses this gap by providing the first zero-knowledge proof system for LoRA fine-tuning of large Transformers, ensuring that the adaptation process is verifiably correct while maintaining privacy.
VIII Conclusion
This work introduces zkLoRA, the first framework to achieve zero-knowledge verifiability for the fine-tuning of large language models using parameter-efficient methods like LoRA. zkLoRA provides end-to-end proofs of correctness for the entire fine-tuning process, including forward and backward propagation as well as parameter updates, without disclosing model parameters or training data. By developing new techniques for handling non-arithmetic operations in Transformer architectures and optimizing the proof system for scalability, zkLoRA enables secure, privacy-preserving model adaptation on billion-parameter models. Experimental results on open-source LLMs demonstrate the practicality and efficiency of the approach, paving the way for trustworthy deployment of fine-tuned LLMs in sensitive or untrusted settings.
Looking ahead, one promising direction is the development of lighter-weight proof systems that further reduce the hardware requirements and prover time. We hope zkLoRA serves as a foundation for future research at the intersection of cryptography and large-scale machine learning, fostering trustworthy deployment of LLMs in security-critical and privacy-sensitive environments.

### v3

@@ V-B Completeness
Statement: VeriLoRA ensures completeness, meaning that a verifier will always accept proofs generated by an honest PPT prover who accurately performs all computations as specified in the VeriLoRA protocol described in Protocol 2.
Analysis: We analyze the completeness of VeriLoRA in its arithmetic computations and non-arithmetic computations, respectively.
  • Arithmetic Operations: VeriLoRA proves the arithmetic computations in LoRA using the sumcheck protocol, which is known to have perfect completeness. This implies that when the prover is honest and performs all computations correctly, the verifier will accept the proof with probability exactly one.
  • Non-arithmetic Operations: VeriLoRA proves the non-arithmetic computations in LoRA by employing lookup-based arguments. Completeness may fail if the random challenge from the verifier (e.g., $\beta$) coincides with a pole of the rational function used in the identity check of lookup-based arguments  [sun2024zkllm]. However, the probability of such an event is bounded by $\frac{C}{|\mathbb{F}|}$, which is negligible for a sufficiently large field size.
In summary, VeriLoRA ensures that the completeness error is negligible, guaranteeing that an honest prover who performs all computations correctly will have an overwhelming probability of convincing the verifier to accept the proof.
@@ V-C Zero Knowledge
Statement: VeriLoRA ensures zero knowledge, meaning that the verifier learns nothing beyond the validity of the statement being proven. Formally, there exists a polynomial-time simulator $\mathcal{S}$ capable of generating transcripts indistinguishable from real interactions, even without access to the prover’s secret witness. The zero-knowledge error is negligible, bounded by $\mathsf{negl}(\lambda)$ under standard cryptographic assumptions.
Analysis: The zero-knowledge property of VeriLoRA is achieved by incorporating the Hyrax-style polynomial commitments [raffel2020exploring] into the sumcheck protocol [sun2024zkllm], which provides computational hiding under the discrete logarithm assumption in the random oracle model. Specifically, VeriLoRA ensures that the verifier learns nothing beyond the validity of the computation while keeping private: 1) the pre-trained model parameters, 2) the low-rank matrices and their updates, 3) the internal computation values in SwiGLU, Softmax, transpose, and element-wise operations during the forward and backward propagations, 4) the training data in the mini-batch.
To rigorously establish the zero-knowledge property of VeriLoRA, we define two executions from the perspective of any PPT adversary $\mathcal{A}$: a real execution with an honest prover and an ideal execution with a simulator. The security objective is to guarantee that the adversary cannot distinguish these two executions.
Formally, VeriLoRA is zero-knowledge if there exists a polynomial-time simulator $\mathcal{S}$ such that, for all PPT adversaries $\mathcal{A}$ and for all witness $w$ (containing pre-trained parameters, LoRA matrices, intermediate computation values, and training data), the following distributions are computationally indistinguishable: $\left|\Pr\left[\mathcal{A}(\mathrm{Real}(w,\mathrm{pp}))=1\right]-\Pr\left[\mathcal{A}(\mathrm{Ideal}(\mathcal{S},\mathrm{pp}))=1\right]\right|\leq\mathsf{negl}(\lambda)$, where $\mathrm{Real}(w,\mathrm{pp})$ and $\mathrm{Ideal}(\mathcal{S},\mathrm{pp})$ denote the real and ideal executions, respectively, and $\mathrm{pp}$ denotes public parameters. The definitions of $\mathrm{Real}$ and $\mathrm{Ideal}$ are given below.
$\mathrm{Real~Execution}~\mathrm{Real}(w,\mathrm{pp})$: 1. $com\leftarrow VeriLoRA\text{-}Commit(w;pp)$, where $w$ includes the base model’s initial weights and the initial low-rank matrices used by LoRA. 2. $\pi\leftarrow VeriLoRA\text{-}Prove(com;pp)$, using the zkMat, zkElementProd, zkTranspose, zkSwiGLU, and zkSoftmax protocols. 3. return $com,\pi$
@@
[…]
In summary, VeriLoRA ensures that the zero-knowledge error is negligible, guaranteeing that the verifier learns nothing beyond the validity of the statement, while preserving the privacy of the prover’s secret witness.
@@ VI Experimental Evaluations
To validate the practical feasibility and performance of VeriLoRA, we conducted comprehensive experiments on real-world large language models and datasets. This section outlines the experimental setup, implementation details, and evaluation results.
@@ VI-A Implementation Details
The implementation of VeriLoRA adopts a hybrid architecture where CUDA components handle core cryptographic operations and zero-knowledge proof generation, while Python scripts provide high-level orchestration and model-specific interfaces. The CUDA modules implement elliptic curve arithmetic and pairing operations on the BLS12-381 curve [boneh2001short], cryptographic commitments for model parameters, and GPU-accelerated zero-knowledge protocols for LoRA Fine-tuning on LLMs. The Python layer manages model downloading, file I/O utilities, and LLaMA-specific processing workflows. We evaluated VeriLoRA for LoRA fine-tuning across six Transformer-based LLMs of varying scales, including LLaMA-3.2 [touvron2023llama] with 3B and 11B parameters, LLaMA-2 [touvron2023llama] with 7B and 13B parameters, and OPT [zhang2022optopenpretrainedtransformer] with 6.7B and 13B parameters, covering models up to 13 billion parameters. We trained the LLMs with the public C4 dataset of English-language text [raffel2020exploring]. Our code is publicly available at https://github.com/liaoguofu/VeriLoRA.
@@ VI-C Rescaling Mechanism for Non-Arithmetic Operations
Figure 2: The proving time of VeriLoRA for different models (in seconds): (a) the forward propagation phase; (b) the backward propagation phase; (c) the parameter update phase, (d) the total.
@@ VI-D1 Proving Time
We evaluated the prover runtime of VeriLoRA across the six Transformer based LLMs. All measurements correspond to the the training process of a mini-batch consists of a single data sample. For each model, we measured the prover runtime of three distinct computational phases of LoRA: forward propagation, backward propagation, and parameter updates. As shown in Fig. 2, the total proving time ranges from 121.93 seconds (LLaMA-3.2-3B) to 249.38 seconds (OPT-13B), reflecting the increasing computational overhead as model size scales. In addition, we measure a LoRA fine-tuning mini-batch (without any ZKPs) executes in 0.47 seconds (LLaMA-3.2-1B), 0.89 seconds (LLaMA-3.2-3B), 1.12 seconds (LLaMA-2-7B), and 1.25 seconds (LLaMA-3.1-8B); relative to these no-zk baselines, our VeriLoRA are approximately three orders of magnitude slower. This order-of-magnitude gap reflects the inherent overhead of zero-knowledge proving in practice. For completeness, we also benchmarked a public zkML inference prover: the distilled GPT-2 inference proof generation takes 7,370.6 seconds [chen2024zkml]. Comparing zkML’s inference proof with our inference prover shows a consistent trend—both are orders of magnitude slower than no-ZK execution—while exact ratios vary with model size and prover stack.
@@ VI-D2 Verification Time
We measured verification time for the six tested LLMs and present the results in Table VI-D4. Across all tested LLMs, verification remains highly efficient. As shown in Table VI-D4, the end-to-end verification time ranges from 1.87 seconds (LLaMA-3.2-3B) to 3.73 seconds (OPT-13B), consistently staying below 4 seconds even for models with over 13 billion parameters. These results demonstrate that the verification workload in VeriLoRA is much more efficient than the proof generation and represents only a minor fraction of the total computational cost.
@@ VI-D3 Polynomial Commitment Cost
We measured the time costs and sizes of the Hyrax polynomial commitments used in VeriLoRA. From the results provided in Table VI-D4, the generation of the commitments introduces a substantial time cost, particularly for large-scale models. The commitment generation time ranges from 156 seconds for LLaMA-3.2-3B to 554 seconds for OPT-13B. This step involves producing polynomial commitments for the different weights of model and dominates the wall-clock latency in our evaluation. The commitment sizes follow a similar trend, increasing from 135.59 MB to 232.67 MB, consistent with the growth in matrix and lookup table dimensionality.
@@ VI-D4 GPU Memory Usage
The total memory usage reported in Table VI-D4 represents the cumulative GPU memory allocated across all Transformer layers during the forward propagation, backward propagation, and parameter update phases while executing VeriLoRA. These memory usages range from approximately 2573 GB for LLaMA-3.2-3B to 3920 GB for LLaMA-2-11B. The computations for the forward propagation, backward propagation, and parameter update phases within the sub-layers of the Transformer layers are executed sequentially on the GPU. Consequently, the actual peak memory observed during VeriLoRA execution remains below 80 GB. This allows VeriLoRA to be efficiently processed on a single NVIDIA A100 GPU with 80 GB of device memory, ensuring practicality even for the largest model scale evaluated.